OpenStudy (cenaida):

Does the data in this table represent linear data? x 1 2 3 4 y 6 12 24 48 A. yes B. no

OpenStudy (cenaida):

@ShaeWX @Tootles143

OpenStudy (anonymous):

Just by first glance, I would say no - but then again it's 2am where I am. 6/1=6 12/2=6 24/3=8 48/4=12 Notice how the values are not equal, that's why I'm suspecting that the data is not linear. Remember that "linear" represents a straight line, and the way I'm looking at that data, the line is straight through the first two points, but after that, it curves. You could also try plotting the points either on a graphing calc or on some good ol' fashioned graph paper and see whether or not the data gives you a linear graph

OpenStudy (cenaida):

Thanks i'll try that!